OpenStudy (abbycross167):

Can someone please help me fix 1 math question? please? I will fan and give a medal? Question - Multiply the following numbers using scientific notation showing the steps as presented in the LIVE recording and in the lesson content. Represent your final answer in both scientific notation and standard notation. (4 x 10^-3)(-2.2 x 10^4).

OpenStudy (abbycross167):

My answer - Step 1 Multiply the number portions. Here, 4 times -2.2 4*-2.2= -8.8 Step 2 Multiply the X10 portions by adding exponents. Here, 10^-3 times 10^4 = 10^-3+4 = 10^1 Answer- (4 x 10^-3)(-2.2 x 10^4) = -8.8 * 10^1 Number in standard notation- Step 1 Move the decimal point 1 place to the right .-.88 Step 2 Check to see if the number makes sense.. 10^1 = 1/100 -8 * 1/100 is -.08. Yes, it makes sense.

OpenStudy (abbycross167):

What my teacher said was wrong - Take your graphing calculator and type in 10^1. does it equal 1/100????

OpenStudy (abbycross167):

@asett do you mind helping me?

OpenStudy (asett):

Sure! Let me read it

OpenStudy (asett):

No, 10^1 does not equal 1/100. Not only is it not a negative exponent, but it also has no real value (the 1) because 10^1 is really just 10. It's that simple.

OpenStudy (asett):

Therefore, all you have to do is move the decimal place to the right in order to get your correct answer, which is -88. Does that make sense? :)

OpenStudy (abbycross167):

yes ma'am it does! thank you again!! you are so helpful :)

OpenStudy (asett):

Oh, good! Just let me know if you need anything else. I'll be here for about another half hour! :)

OpenStudy (abbycross167):

yes ma'am thank you :)